Ticket Number: HEA-NOC/20100606-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20100606 10:27 UTC+1 Problem Start: 20100606 09:40 UTC+1
Ticket Update: 20100608 10:19 UTC+1 Problem End:
Site/line: NUIM, Schools Network, Wicklow
Outage and flapping on circuit to nuim-sw3.
Loss of resilience to NUIM, due to loss of connectivity on backup link
Loss of connectivity to:
IT Carlow Wicklow campus.
NUIM Hamilton - CERN P2P
NUIM Hamilton - WIT P2P
Schools High Speed: BL - NUIM 1 & 2, CWT - NUIM 1 & 2, Kilkenny - NUIM
20100606 10:12 CS Reported to UPC, reference 440320
20100606 12:42 CS Called UPC, no update on ticket, call centre will call
the technical staff and ask for an update.
20100606 14:13 CS No word from UPC, called them again. The circuit number
had been misheard/mistyped. Gave them correct circuit number. They will
pass the ticket to the technical support and they will call back.
20100606 14:26 CS Got call back to confirm A & B ends, they are
20100606 15:29 CS Called UPC engineer for update, they've checked as much
as they can remotely and need site access at NUIM to check the fibre.
20100606 15:59 CS Called Liam Burbridge to verify no site problems and to
arrange access for UPC. Gave him name & number of UPC engineer.
20100606 16:05 CS Updated list of affected clients.
20100606 17:37 CS UPC were onsite at NUIM and saw no problems with the
fibre - RX power from Blanch was fine, transmit power was fine.
LambdaDriver unit in NUIM had no error lights. Have logged onto Blanch unit
and tried a cold reset, to no avail. Can't obtain any useful information
from unit as to status of fibre. It is possible there is a uni-directional
fibre break, but this should be detected by the bluenet switches and has
20100606 21:12 BN Further investigation with the help of NUIM shows that
the local optics and fibre are working. Bluenet switch appears to be fine,
suggesting that the issue lies with the intervening HEAnet optical
equipment. Further investigation and consideration of plans for resumption
of service are ongoing.
20100607 22:24 GAR Due to unavailability of spares we have been unable to
restore this circuit to date. All NUIM services remain failed over to the
resilient link but the Layer 2 circuits remain out of service.
An alternate method of restoring the link has been identified and work on
commissioning this will begin first thing Tuesday morning.
20100608 10:17 GAR On site in NUIM and Blanchardstown, the A- and B-ends
of the failed link. Commencing reseating of cards and hard power cycling of
Lambda Driver equipment in an effort to restore connectivity.
This ticket can be monitored at http://www.hea.net/tickets/20100606-1
HEAnet Limited. Registered in Ireland, No. 275301.